OpenStudy (1prettylisa):

How to simplify ³√16

OpenStudy (narissa):

Step 1-Rewrite 16 as its prime factor. ³√2x2x2x2 Step 2-Group the same prime factors into groups of three ³√(2x2x2)x2 Step 3-Rewrite each group of three in exponent form ³√2^3x2 Step 4- Apply this rule;³√x^3=x Anwer---2³√2 Or the answer can also be... Step 5-Use product rule x^a x^b=x^a+b Answer---2 4/3

OpenStudy (3mar):

\[\sqrt[3]{16}=\sqrt[3]{8*2}=\sqrt[3]{8}*\sqrt[3]{2}=2\sqrt[3]{2}\]

OpenStudy (narissa):

Sorry it took me a while to explain...

OpenStudy (3mar):

as \[8=2^3\] \[\sqrt[3]{8}=2\]